OpenStudy (anonymous):

Curium-243 has a half-life of 28.5 days. In a sample of 5.6 grams of curium-243, how many grams will remain after 12 days?

OpenStudy (anonymous):

There are two parts to this problem. Here's how to do the first.\[y=Ce^{kt},\text{ where}\\ y=\text{amount of substance after time t}\\ C=\text{initial amount of substance}\\ k=\text{relative decay factor}\\ t=\text{time}\] You're given that Curium-243 has a half-life of 28.5 days, so after this period of time, you have \[\frac{1}{2}=e^{28.5k}\]And solve for k.

OpenStudy (anonymous):

Second part: Using the k you found in the first part, use the next set of information and the formula: \[y=5.6e^{12k}\] And solve for y.

OpenStudy (anonymous):

i got a really big number i dk what i did wrong

OpenStudy (anonymous):

\[\frac{1}{2}=e^{28.5k}\\ \ln\left(\frac{1}{2}\right)=28.5k\\ k=\frac{\ln\left(\frac{1}{2}\right)}{28.5}\\ k\approx-0.0243\] Did you get this?